# Biomechanics-Aware Trajectory Optimization for Navigation during Robotic Physiotherapy

### v1

@@ II-A1 Generating differentiable strain maps
where $l_{T}(\boldsymbol{x},\alpha)$ is the current length of the tendon and $l_{0}$ its slack length at rest. Querying the model for tendon strains during motion planning is impractical due to the computationally intensive calculations involved. Similar to our previous work, we can fully describe the change in strain $\sigma$ based on the human model’s state via pre-computed “strain maps” [30, 31, 38]. In the present work, we ignore the dependence on tendon velocity since rehabilitation movements are generally slow. With this assumption, we spanned the space of possible arm configurations and rotator cuff muscle activations to query the model for the tendon strain induced in every tendon of the rotator cuff. During this pre-computation phase, fixed 4-degree steps were used to vary the movement DoFs within the shoulder range of motion, while muscle activation was varied at 0.005 discretization. In this way, we created high-dimensional strain maps for every individual tendon, which allowed us to target only one tendon at a time or, alternatively, to monitor the strain on the rotator cuff as a whole by constructing aggregate maps that consider the highest strain across all the tendons [30, 38]. To simplify the understanding of strain level for non-expert users of our system (e.g., physiotherapists and patients), we chart $\sigma$ on 2D grids where the strain percentage is visualized as a function of $PE$ and $SE$ angles, while both $AR$ and the muscular activation $\alpha$ are fixed. To consider different values of $AR$ and $\alpha$, other (pre-computed) 2D maps can be traversed.
@@
As a final step, we interpolated our discrete strain maps using 2D Gaussian radial basis functions to leverage more efficient gradient-based methods in our trajectory optimization problem (Section II-B). Through least-squares fitting, we obtain the parameter vector $\boldsymbol{\pi}\in\mathbb{R}^{N_{p}\times N_{G}}$ that describes the combination of the $N_{G}$ Gaussian functions best approximating every 2D strain map, in terms of the $N_{p}$ parameters of every Gaussian.
II-A2 Obtaining differentiable skeletal dynamics
@@
after computing navigable differentiable maps of rotator cuff tendon strain, we need to account for the human movement dynamics governing how the human shoulder should be moved by the robot. This information can be obtained from the OpenSim model only numerically, and this has led researchers to employ less efficient gradient-free optimization methods to optimize human-robot interaction [26]. To enhance computational efficiency, we employed OpenSimAD [39] as a tool that allows us to retrieve a differentiable function from an OpenSim musculoskeletal model. OpenSimAD is an instrumented version of OpenSim recently developed to automatically generate the differentiable version of OpenSim outcomes and functions. It records the entire call sequence of the function used to compute a particular outcome variable of interest within the OpenSim model, offering the possibility to employ this call sequence for algorithmic differentiation by programs like CasADi[40]. In particular, we customized the OpenSimAD pipeline to obtain a differentiable representation of our reduced-order model of the human shoulder in terms of a set of ordinary differential equations that describe the dynamics of the state $\boldsymbol{x}\in\mathbb{R}^{N_{x}}$ of the multibody skeletal system, given a set of generalized forces $\boldsymbol{u}\in\mathbb{R}^{N_{q}}$ acting on the model’s DoFs:
$\dot{\boldsymbol{x}}=f(\boldsymbol{x},\boldsymbol{u}),$ (3)
@@
where $N_{q}=3$ and $N_{x}=2N_{q}$ in our case. Importantly, this procedure guarantees that the model’s joint definitions will be preserved, together with personalized dimensions, mass, and inertia for every existing bone in the original model.
@@ II-B Biomechanics-aware trajectory optimization
We combine information from the differentiable strain maps and human dynamics to formulate a biomechanics-aware optimal control problem (OCP), with the goal of synthesizing safe human motion towards a target shoulder position. The solution to the OCP is the optimal trajectory $\boldsymbol{x}^{*}$ and corresponding generalized forces $\boldsymbol{u}^{*}$ driving the human model toward the goal pose. The sequence of optimal controls and states $\{\boldsymbol{x}^{*},\boldsymbol{u}^{*}\}$ is then transformed into a reference for the robot controller to deliver optimal rehabilitation motion to the human subject.
We opted to plan trajectories that are fully contained in a single 2D strain map, so that the output of our optimization would be intuitive for non-expert users and easy to visualize in two dimensions (e.g., on a computer screen). We selected plane of elevation ($PE$) and shoulder elevation ($SE$) to be our controlled DoFs, while current values of axial rotation ($AR$) and muscle activation $\alpha$ are input to select the strain map to plan onto.
We define below the key elements of the OCP (i.e., cost function and constraints) that allow us to set the optimality criterion with respect to which $\{\boldsymbol{x}^{*},\boldsymbol{u}^{*}\}$ are found.
@@ II-B1 Cost Function
we design our cost function to capture important requirements for a safe rehabilitation trajectory: the movement should minimize the strain on selected tendons, produce low accelerations on the human body, and evolve towards the target final position. These elements are mirrored into the definition of our cost $L(\boldsymbol{x}(t),\dot{\boldsymbol{x}}(t))$, formalized as a weighted sum of the following terms:
    $L_{\sigma}=\sigma(\boldsymbol{x},\alpha)$, accounting for the instantaneous (non-negative) strain that the selected tendon undergoes;
    $L_{\text{acc}}=\dot{\boldsymbol{x}}^{T}\boldsymbol{Q_{1}}\dot{\boldsymbol{x}}$, accounting for the derivatives of our state variables during the motion. The positive semidefinite matrix $\boldsymbol{Q_{1}}\in\mathbb{R}^{6\times 6}$ allows us to consider only the relevant accelerations when planning on the current strain map;
    $L_{\text{$\tau$}}=\frac{1}{d_{0}^{2}}(\boldsymbol{x}-\boldsymbol{x}_{\text{$\tau$}})^{T}\boldsymbol{Q_{2}}(\boldsymbol{x}-\boldsymbol{x}_{\text{$\tau$}})$, accounting for the distance to the target human model pose $\boldsymbol{x}_{\tau}$, where $d_{0}$ represents the distance to the target pose at the beginning of the trajectory, to normalize the contribution of this term. Again, $\boldsymbol{Q_{2}}\in\mathbb{R}^{6\times 6}$ is a positive semidefinite matrix selecting only the relevant human DoFs.
In the following, we will refer to the weights of these terms as $w_{\sigma}$, $w_{\text{acc}}$ and $w_{\tau}$ respectively (where the latter does not include the normalization by $d_{0}$ introduced above).
@@ II-B2 Constraints
    Initial condition: every new trajectory $\{\boldsymbol{x}^{*},\boldsymbol{u}^{*}\}$ should start from the current state of the human model $\hat{\boldsymbol{x}}_{\text{curr}}$, estimated through the robot’s encoders (see Sec. II-D). The current state is used as the initial condition for the next instance of the OCP enforcing $\boldsymbol{x}(t=0)=\hat{\boldsymbol{x}}_{\text{curr}}$, where the initial time is set to be 0 without loss of generality.
    Torque limits: to make sure that following the optimal trajectory does not require excessive torque to be exerted along the human DoF, we limit them within heuristic bounds by requiring $\boldsymbol{u}_{\text{min}}\leq\boldsymbol{u}(t)\leq\boldsymbol{u}_{\text{max}}$, where $\boldsymbol{u}_{\text{min}},\boldsymbol{u}_{\text{max}}\in\mathbb{R}^{3}$. Bounds can be adjusted to enforce different toque limits on the various DoF accounting, for example, for the fact that torques along SE should counteract gravity.
    Terminal condition: to impose an acceptable final state for the human model, we prescribe that $(\boldsymbol{x}_{t=T_{f}}-\boldsymbol{x}_{\tau})\odot(\boldsymbol{x}_{t=T_{f}}-\boldsymbol{x}_{\tau})\leq\boldsymbol{\epsilon}$, where $T_{f}$ represents the length of the planning horizon, and $\odot$ the Hadamard product. If reaching the target final state is possible in the current OCP instance (e.g. if $T_{f}$ is long enough), the parameter $\boldsymbol{\epsilon}\in\mathbb{R}^{6}$ is defined for every element of the state vector, meaning both final pose and velocity of the human model can be specified fully. Otherwise, if the target pose specified by $\boldsymbol{x}_{\tau}$ is too far from $\hat{\boldsymbol{x}}_{\text{curr}}$, only the final velocities are limited. This choice guarantees low velocities of the human arm at the end of the planning horizon and ensures recursive feasibility through a safe terminal set $\mathbb{S}_{T}$.
II-C Optimal control problem (OCP) transcription
@@
$\quad\boldsymbol{x}(0)-\hat{\boldsymbol{x}}_{\text{curr}}=\boldsymbol{0}$ initial state
@@
$\quad\boldsymbol{x}_{t=T_{f}}\in\mathbb{S}_{T}$ terminal set
@@
To solve the OCP, we cast it into discrete time, transcribing (4) into an equivalent Non-Linear Programming problem (NLP) that can be solved by structure-exploiting solvers. We achieved this with orthogonal collocation techniques [41], hence approximating the state trajectories with suitable dth-order polynomial splines. The overall optimization horizon $T_{f}$ is broken down into $N$ intervals of equal length, and inside the generic interval $[t_{k},t_{k+1}]$ we select $d=3$ Legendre-Gauss collocation points at which the dynamics as in (3) are enforced. This step results in additional continuity and collocation constraints that need to be considered in the NLP formulation.
@@
We implemented the resulting NLP in Python and solved it with IPOPT [42] leveraging MA27 [43] as a linear solver, while first- and second-order derivatives were provided by CasADi [40] and OpenSimAD. To clarify the computational advantage of using the differentiable skeletal dynamics, we also embedded numerically evaluated gradients with perturbation calls to the original OpenSim model directly in the collocation constraints of the NLP, through CasADi. We ran simulation experiments for both cases, to compare the computational requirements for both choices.
II-D Human state estimation
@@
Figure 3: Overview of the coordinate systems of the setup: The shoulder frame (green) originates in the center of the glenohumeral joint. The glenohumeral joint DoFs (PE, SE, and AR) are shown in red. The elbow frame (blue) originates in the center of the elbow, and a fixed transformation relates it to the robot’s end-effector frame (pink).
@@
II-E Robot control and model-based gravity compensation
Apart from being used for human pose estimation, the key purpose of the robotic arm is to deliver therapeutic motion to the human subject based on the continuously optimized trajectory. We map the optimal trajectory $\boldsymbol{x}^{*}$ resulting from (4) into the corresponding robot’s end-effector reference pose:
${}^{\text{base}}\bar{\boldsymbol{R}}_{\text{EE}}={}^{\text{base}}\boldsymbol{R}_{\text{sh}}{}^{\text{sh}}\boldsymbol{R}_{\text{elb}}(\boldsymbol{x}^{*}){}^{\text{elb}}\boldsymbol{R}_{\text{EE}}$ (11)

### v2

@@ II-A1 Navigable strain maps
where $l_{i}(\boldsymbol{x})$ is the current length of the tendon and $l_{0,i}$ its slack length at rest. We fully describe the change in strain $\sigma_{i}$ based on the human model’s state via pre-computed “strain maps” [38, 39, 45], for every tendon of interest. In this way, we created high-dimensional strain maps for every individual tendon, which allowed us to target only one tendon at a time or, alternatively, to monitor the strain on the rotator cuff as a whole by constructing aggregate maps that consider the highest strain across all the tendons [38, 45]. To simplify the understanding of strain level for non-expert users of our system (e.g., physiotherapists and patients), we chart $\sigma_{i}$ on 2D grids where the strain percentage is visualized as a function of $PE$ and $SE$ angles, while both $AR$ and $\alpha_{i}$ are fixed. To consider different values of $AR$ and $\alpha_{i}$, other (pre-computed) 2D maps can be traversed.
@@
To traverse over strain maps efficiently, we interpolated our discrete strain maps using 2D Gaussian radial basis functions. Through least-squares fitting, we obtain the parameter vector $\boldsymbol{\pi}\in\mathbb{R}^{N_{p}\times N_{G}}$ that describes the combination of the $N_{G}$ Gaussian functions best approximating every 2D strain map, in terms of the $N_{p}$ parameters of every Gaussian.
II-A2 Controlling skeletal dynamics
@@
After computing navigable differentiable maps of rotator cuff tendon strain, we account for the human movement dynamics when the human shoulder is being manipulated by the robot. Human dynamics should be used to inform the trajectory planner (Sec. II-B) about the relationship between forces/torques that the robot applies to the human, and the resulting human accelerations. However, the OpenSim API provides this relationship only in a non-differentiable form, which led previous research [31] to employ less efficient gradient-free optimization methods to optimize human-robot interaction. To efficiently represent human skeletal dynamics for control purposes, we employed OpenSimAD [46] as a tool that generates differentiable functions from OpenSim musculoskeletal models. OpenSimAD extends OpenSim by recording the call sequence used to compute specific outcome variables within an OpenSim model, enabling algorithmic differentiation through frameworks like CasADi [47]. In particular, we customized the OpenSimAD pipeline to obtain a differentiable representation of our human model in terms of the ordinary differential equations capturing the dynamics of its multibody skeletal system, given a set of generalized forces $\boldsymbol{u}\in\mathbb{R}^{N}$ and the model’s kinematic state $\boldsymbol{\mathrm{q}}\in\mathbb{R}^{N_{q}}$:
$\dot{\boldsymbol{\mathrm{q}}}=f(\boldsymbol{\mathrm{q}},\boldsymbol{u}),$ (5)
@@
where $N=3$ and $N_{q}=2N$ in our case. Importantly, this procedure guarantees that the model’s joint definitions will be preserved, together with personalized dimensions, mass, and inertia for every existing bone in the original model.
@@ II-B Biomechanics-aware trajectory optimization
We combine information from the differentiable strain maps and human skeletal dynamics to formulate a biomechanics-aware optimal control problem (OCP), with the goal of synthesizing safe human motion towards a target shoulder position. The solution to the OCP is the optimal trajectory $\boldsymbol{\mathrm{q}}^{*}$ and corresponding generalized forces $\boldsymbol{u}^{*}$ driving the human model toward the goal pose. Note that we optimize generalized forces rather than end-effector wrenches to keep our trajectory optimization agnostic to how rehabilitation trajectories are delivered to the patient. The actual reference for the robot controller is obtained from the optimal sequence of kinematic states and controls $\{\boldsymbol{\mathrm{q}}^{*},\boldsymbol{u}^{*}\}$ in Sec. IV-A.
We opted to plan trajectories that are fully contained in a single 2D strain map, so that the output of our optimization would be intuitive for non-expert users and easy to visualize in two dimensions (e.g., on a computer screen). We selected plane of elevation ($PE$) and shoulder elevation ($SE$) to be our controlled DoFs, while current values of axial rotation ($AR$) and muscle activation $\boldsymbol{\alpha}$ are input for trajectory planning to select the correct strain map.
We define below the cost function and constraints of the OCP that allow us to set the optimality criterion with respect to which $\{\boldsymbol{\mathrm{q}}^{*},\boldsymbol{u}^{*}\}$ are found, and present how the OCP is solved in practice.
@@ II-B1 Cost Function
we design our cost function to capture important requirements for a safe rehabilitation trajectory: the movement should minimize the strain on selected tendons, produce low accelerations on the human body to reduce interaction forces and discomfort, and evolve towards the target final position. These elements are mirrored into the definition of our cost $L(\boldsymbol{x},\boldsymbol{u})$, formalized as a weighted sum of the following terms:
  • $L_{\sigma}=\sigma(\boldsymbol{x})$, accounting for the instantaneous (non-negative) strain that the selected tendon group undergoes;
  • $L_{\text{acc}}=\dot{\boldsymbol{\mathrm{q}}}^{T}\boldsymbol{Q_{1}}\dot{\boldsymbol{\mathrm{q}}}$, accounting for the derivatives of our relevant state variables during the motion. The positive semidefinite matrix $\boldsymbol{Q_{1}}\in\mathbb{R}^{6\times 6}$ selects only joint accelerations when planning on the current strain map;
  • $L_{\text{$\tau$}}=\frac{1}{d_{0}^{2}}(\boldsymbol{\mathrm{q}}-\boldsymbol{\mathrm{q}}_{\text{$\tau$}})^{T}\boldsymbol{Q_{2}}(\boldsymbol{\mathrm{q}}-\boldsymbol{\mathrm{q}}_{\text{$\tau$}})$, accounting for the distance to the target human model pose $\boldsymbol{\mathrm{q}}_{\tau}$, where $d_{0}$ represents the distance to the target pose at the beginning of the trajectory, to normalize the contribution of this term. Again, $\boldsymbol{Q_{2}}\in\mathbb{R}^{6\times 6}$ is a positive semidefinite matrix selecting only the relevant human DoFs.
Overall, the cost to be minimized is:
$L(\boldsymbol{x},\boldsymbol{u})=w_{\sigma}\sigma(\boldsymbol{x})+w_{\text{acc}}\dot{\boldsymbol{\mathrm{q}}}^{T}\boldsymbol{Q_{1}}\dot{\boldsymbol{\mathrm{q}}}+w_{\tau}\frac{1}{d_{0}^{2}}\boldsymbol{\Delta}_{\mathrm{q}}^{T}\boldsymbol{Q_{2}}\boldsymbol{\Delta}_{\mathrm{q}}$ (6)
where $w_{\sigma}$, $w_{\text{acc}}$ and $w_{\tau}$ are scalars weighting the contribution of the various terms, and we indicated $(\boldsymbol{\mathrm{q}}-\boldsymbol{\mathrm{q}}_{\text{$\tau$}})=\boldsymbol{\Delta}_{\mathrm{q}}$ for brevity.
@@ II-B2 Constraints
    Initial conditions: every new trajectory $\{\boldsymbol{\mathrm{q}}^{*},\boldsymbol{u}^{*}\}$ should start from the current kinematic state of the human model $\hat{\boldsymbol{\mathrm{q}}}_{\text{curr}}$, estimated through the robot’s encoders (see Sec. III-B). This value is used as the initial condition for the next instance of the OCP enforcing $\boldsymbol{\mathrm{q}}(t=0)=\hat{\boldsymbol{\mathrm{q}}}_{\text{curr}}$, where the initial time is set to be 0 without loss of generality.
    Torque limits: to ensure that following the optimal trajectory does not require excessive torque to be exerted along the human DoF, we limit them within heuristic bounds by requiring $\boldsymbol{u}_{\text{min}}\leq\boldsymbol{u}(t)\leq\boldsymbol{u}_{\text{max}}$, where $\boldsymbol{u}_{\text{min}},\boldsymbol{u}_{\text{max}}\in\mathbb{R}^{3}$. Bounds can be adjusted to enforce different torque limits on the various DoF accounting, for example, for the fact that torques along SE should counteract gravity.
    Terminal condition: to impose an acceptable final state for the human model, we prescribe that $(\boldsymbol{\mathrm{q}}_{t=T_{f}}-\boldsymbol{\mathrm{q}}_{\tau})\odot(\boldsymbol{\mathrm{q}}_{t=T_{f}}-\boldsymbol{\mathrm{q}}_{\tau})\leq\boldsymbol{\epsilon}$, where $T_{f}$ represents the length of the planning horizon, and $\odot$ the Hadamard product. If reaching the target final state is possible in the current OCP instance (e.g. if $T_{f}$ is long enough), the parameter $\boldsymbol{\epsilon}\in\mathbb{R}^{6}$ is defined for every element of $\boldsymbol{\mathrm{q}}$, meaning both final pose and velocity of the human model can be specified fully. Otherwise, if the target pose specified by $\boldsymbol{\mathrm{q}}_{\tau}$ is too far from $\hat{\boldsymbol{\mathrm{q}}}_{\text{curr}}$, only the final velocities are limited. This choice guarantees low velocities of the human arm at the end of the planning horizon and ensures recursive feasibility through a safe terminal set $\mathbb{S}_{T}$ when the optimization is performed iteratively over a receding time horizon.
II-B3 Optimal control problem (OCP) transcription
@@
$\quad\boldsymbol{\mathrm{q}}(0)-\hat{\boldsymbol{\mathrm{q}}}_{\text{curr}}=\boldsymbol{0}$ initial state
@@
$\quad\boldsymbol{\mathrm{q}}_{t=T_{f}}\in\mathbb{S}_{T}$ terminal set
@@
To solve the OCP, we cast it into discrete time, transcribing (7) into an equivalent Non-Linear Programming problem (NLP) that can be solved by structure-exploiting solvers. We used orthogonal collocation techniques [48], hence approximating the state trajectories with suitable dth-order polynomial splines. The overall optimization horizon $T_{f}$ is broken down into $N$ intervals of equal length, and inside the generic interval $[t_{k},t_{k+1}]$ we select $d=3$ Legendre-Gauss collocation points at which the dynamics as in (5) are enforced. This step results in additional continuity and collocation constraints that need to be considered in the NLP formulation. Resulting optimal trajectories $\{\boldsymbol{\mathrm{q}}^{*},\boldsymbol{u}^{*}\}$ were upsampled for smoother execution.
@@
We implemented the resulting NLP in Python and solved it with IPOPT [49] leveraging MA27 [50] as a linear solver, while derivatives were provided by CasADi [47] and OpenSimAD. To clarify the computational advantage of using the differentiable skeletal dynamics, we also embedded numerically evaluated gradients with perturbation calls to the original OpenSim model directly in the collocation constraints of the NLP, through CasADi. We ran simulations for both cases to compare computational requirements.
III Dynamic Adaptation to Human Actions
The trajectory on strain maps generated by the planner (Sect. II-B) provides a reference for the robot controller. Nevertheless, unpredictable volitional actions of the patient and/or reflexive reactions to physical human-robot interaction can induce large interaction forces and corresponding muscle activations that lead to potentially large changes in the strain map. In this section, we describe how we exploit the robot’s sensory system to make real-time updates on estimates of patient pose and muscle activity. These estimates are used to update the strain maps, which the planner will use to re-plan the robot trajectories in an online manner to avoid dangerous high-strain areas and abide by given optimization objectives.
III-A Online estimation of muscle activity
Estimating the activity of deep rotator cuff muscles is not feasible using traditional surface electromyography (EMG) measurements. To overcome this limitation, we employed a model-based estimation approach to infer individual muscle activations within our biomechanical model during robot-assisted rehabilitation. Specifically, we used the Rapid Muscle Redundancy (RMR) solver [27, 45], which yields physiologically plausible muscle activations in over-actuated systems — i.e., systems with more muscles than degrees of freedom — while accounting for the stabilizing role of the rotator cuff at the glenohumeral joint [27]. As the human musculoskeletal model is inherently redundant, the solver leverages nonlinear optimization to estimate muscle activations consistent with observed kinematics and forces. The RMR solver minimizes a muscle effort cost function subject to physiological constraints to generate realistic recruitment patterns.
The original open-source implementation of RMR was designed for offline use. In contrast, our application requires real-time estimation, with the model state updated continuously based on measured joint kinematics (Section III-B) and external human-robot interaction wrenches (Section IV-B). To this end, we restructured the solver into a class-based module integrated with the Robot Operating System (ROS), enabling online estimation of muscle activations $\hat{\boldsymbol{\alpha}}$ in a dedicated sub-process. These real-time estimates were used to select the tendon strain map that informed the trajectory optimization described in Sec. II-B.
The code implementation is available in the project repository at https://anonymous.4open.science/r/baton-robotic-rehab-2545.
III-B Human kinematics estimation
@@
Figure 3: The human and robot coordinate systems and their correspondence. The shoulder frame (green) originates in the center of the glenohumeral joint. The glenohumeral joint DoFs (PE, SE, and AR) are shown in red. The elbow frame (blue) originates in the center of the elbow, and a fixed transformation relates it to the robot’s end-effector frame (pink).
We estimated the human model’s kinematic state $\hat{\boldsymbol{\mathrm{q}}}$ through the position, orientation, and twist of the robot’s end-effector (EE), which we mapped to the shoulder state $\boldsymbol{x}$ according to the fixed rigid connection between the EE and elbow locked at $90^{\circ}$(see Fig. 3). An estimation of the full human body pose is out of the scope of our current work, so we assume for simplicity that the center of the glenohumeral joint is stationary during the experiments and that the orientation of the human torso is fixed. Under these conditions, monitored during the experiments, it is possible to reconstruct the shoulder kinematic state through the robot’s encoders.
@@
Finally, we differentiated numerically $\hat{\dot{PE}}$, $\hat{\dot{SE}}$, and $\hat{\dot{AR}}$ to obtain the corresponding acceleration vector $\boldsymbol{\ddot{\theta}}$. Exponential filtering was applied to the estimated values to reject high-frequency components incompatible with human movements in physiotherapy.
IV Robot control and experimental design
To test BATON as a novel approach to robotic-assisted rotator cuff rehabilitation and demonstrate its capabilities, we designed and performed various lab-based experiments. These proof-of-concept scenarios are essential for the validation of our approach as a whole, testing BATON against the state of the art and in situations where its dynamic adaptation becomes crucial for guaranteeing safe rehabilitation.
IV-A Robot control
In addition to human state estimation, our collaborative robotic arm administers therapeutic motion to the human subject. We map the optimal trajectory $\boldsymbol{\mathrm{q}}^{*}$ resulting from (7) into the corresponding robot’s end-effector reference pose:
${}^{\text{base}}\bar{\boldsymbol{R}}_{\text{EE}}={}^{\text{base}}\boldsymbol{R}_{\text{sh}}{}^{\text{sh}}\boldsymbol{R}_{\text{elb}}(\boldsymbol{\mathrm{q}}^{*}){}^{\text{elb}}\boldsymbol{R}_{\text{EE}}$ (14)